Need some help with this one.
For the funtion F(x)=X^2-2x+1
A) Find F(0)
B) Solve F(x)=0

Answer by stanbon(75887)

1. F(0) means you should put "0" in place of "x".
F(0)=0^2-2(0)+1 = 1
2. If F(x)=0, the x^2-2x+1=0
(x-1)^2=0
Therefore x= 1 and x=1 (1 is a double-root)
Cheers,
stan H.